OpenStudy (anonymous):

For the following series circuit, compute the power in R1, R2, and R3 if the circuit values are E = 24 volts, R1 = 8, R2= 12, and R3 = 28.

OpenStudy (anonymous):

Hi. Long time since I did anything like this, but see if this looks OK... Resistances are in series, so total resistance of circuit, Rtot, is given by Rtot=R1+R2+R3 Therefore, Rtot=48 ohm Current in circuit, I, is given by I=V/R, where V=voltage, R=resistance Therefore I=0.5 amp P=Isquared R, where P=power Therefore, in R1, P=2 watt; in R2, P=3 watt; and in R3, P=7 watt My first post - only found this site a few minutes ago. Next time I'll use the equation-writing tool. Best regards.

OpenStudy (anonymous):

Thank you!! This section is killing me so any help whatsoever is appreciated.

OpenStudy (anonymous):

Happy to help. I was great at UK school physics to age 16, lousy at it from 16 to 18. Fortunately, I did circuit stuff when I was 15.

OpenStudy (anonymous):

Wow, you did this stuff when you were young, and you had to do it for three years? I'm dying with just one.

OpenStudy (anonymous):

Did two years at what was then called O-level (O for ordinary), and then two years at what's still called A-level (A for advanced). Ouch to the second one.

OpenStudy (anonymous):

Physics wasn't compulsory at either stage. My pick x2.

OpenStudy (anonymous):

Ahh cool. Unfortunately my school has very limited options for classes so we don't have anything like that.